The DRI for protein is about 0.80 g/kg body weight. How much protein is required each day by a 210-lb football player? Express your answer to two significant figures and include the appropriate units.
Answer #1

weight of football player is 210 lb converting it to kg

1 lb =0.45359 kg so 220 lb= 210*0.45359= 95.25 kg

DRI of the protein is 0.80 g/kg so

1 kg of weight requires 0.80 g

99.79 kg of weight will require = 95.25*0.80= 76.2 g = 76  grams